OpenStudy (anonymous):

5 minus -2-6+8=9 right

OpenStudy (radar):

right, 5+2-6+8=15-6=9

OpenStudy (anonymous):

No, 5.

OpenStudy (anonymous):

You fooled me; I thought the minus minus was a typo..